Circular No. 8189

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SATELLITES OF MARS D. K. Lynch, R. W. Russell, D. L. Kim, H. B. Hammel, and R. B. Perry

report that 3- to 13-μm spectroscopy of the satellites of Mars on Aug. 21.6 UT, obtained as noted for V2573 Oph on IAUC 8186, showed significant departures from a single-temperature gray-body function. The 5- to 12-μm color temperatures of Mars II (Deimos) and Mars I (Phobos) were 310 ± 5 and 345 ± 5 K, respectively — significantly higher than the blackbody thermal equilibrium temperature (237 K) at Mars’ heliocentric distance. Relative to the color-temperature fits, Deimos exhibited significant excess emission between 8 and 10.5 μm, and Phobos showed only marginal excess emission. These findings are based on a flux model for the calibrator α Tau that includes the SiO absorption between 7.9 and 10 μm; similar results were found when the data were reduced relative to α Lyr. Measured infrared magnitudes: Phobos, L = 6.3 ± 0.07, M = 4.2 ± 0.07, [N(10.2 μm)] = 0.05 ± 0.07; Deimos, M = 5.3 ± 0.3, [N(10.2 μm)] = 0.76 ± 0.05.

SUPERNOVAE 2003gv, gw, hc, he, hf, hi, hj, hk, hl, hm, 2003hp A. V. Filippenko and R. J. Foley, University of California, Berkeley;

and F. J. D. Serduke, Lawrence Livermore National Laboratory, report that inspection of CCD spectra (range 320–1000 nm), obtained on Aug. 28 UT with the Shane 3-m telescope at Lick Observatory, reveals that SN 2003hj (IAUC 8184) is of type Ia, ∼ 2 weeks past maximum brightness, at redshift 0.076 (determined from narrow Hα emission in the host galaxy nucleus). Supernovae 2003hi (IAUC 8184), 2003hl (IAUC 8184), 2003hk (IAUC 8184), 2003gv (IAUC 8175), and 2003hc (IAUC 8179) are all relatively normal type-II supernovae, with hydrogen Balmer lines having P-Cyg profiles. SN 2003gw (IAUC 8176) is also probably of type II, but the noisy spectrum exhibits only rather narrow (FWHM = 2000 km/s) Hα emission. SN 2003hm (IAUC 8185) is of type Ia, ∼ 2 months past maximum brightness. Supernovae 2003he (IAUC 8182) and 2003hp (IAUC 8188) show some broad undulations and are probably supernovae at respective redshifts of 0.025 and 0.051 (determined from narrow Hα emission in the host galaxy nuclei), though of uncertain types because their spectra are noisy; the former has a very red continuum, while the latter is reasonably blue. SN 2003hf (IAUC 8182) might be of type II, exhibiting a featureless blue continuum, but such characteristics are rare for a type-II supernova as old as SN 2003hf (at least 2 weeks past the explosion).

2003 August 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green